hallm:

Which function corresponds to the table? A) y = 3x - 2 B) y = 2x + 3 C) y = -2x + 3 D) y = -3x + 2

queenmel:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hallm Which function corresponds to the table? A) y = 3x - 2 B) y = 2x + 3 C) y = -2x + 3 D) y = -3x + 2 \(\color{#0cbb34}{\text{End of Quote}}\) what table

hallm:



hallm:

that one and the first y is a 3 i know it looks weird im sorry

queenmel:

is it a)

hallm:

thank you

queenmel:

welcome